OpenStudy (liv33):

Fan & Medal... What is the value of -2x^3+5x-7, when x=-3 A. 22 B. 32 C. 36 D. 46

OpenStudy (liv33):

D is supposed to be 40 not 46 sorry

OpenStudy (greenbannanabob):

Just plug in -3 in place of x. -2(-3)^3+5(-3)-7=?

OpenStudy (liv33):

so... 6^3+2-7= 216+(-5)= 211 ?

OpenStudy (liv33):

That's what I did before, but that isn't an option

OpenStudy (greenbannanabob):

Not exactly. -2(-3)^3 isn't 216. You first do -3^3 and then multiply by -2.

OpenStudy (liv33):

so -27(-2)

OpenStudy (liv33):

which is 54

OpenStudy (greenbannanabob):

Yes

OpenStudy (liv33):

is that it? it isn't an answer

OpenStudy (greenbannanabob):

No, it's 54+(-5)

OpenStudy (liv33):

ok so that's -10.8

OpenStudy (greenbannanabob):

Wait, you made another mistake 5(-3) doesn't equal 2. It means 5*-3 which is -15. That means that the equation was supposed to be 54-15-7

OpenStudy (greenbannanabob):

Simplify that and you'll have your answer

OpenStudy (liv33):

32

OpenStudy (liv33):

Thanks a lot :)

OpenStudy (greenbannanabob):

You're welcome